Question 1 (1 point) Below is the balance sheet for a bank. Under Other it has listed $X just think of this as the dollar
Answer #1

The required reserve ratio is 6%. Thus, the required reserve amount will be 6%*(Total Liability)=0.06*(Total Asset)=0.06*(36+160+44)=0.06*240=14.4

Thus, the level of required reserve is 14.4(Ans)